OpenStudy (anonymous):

What are the explicit equation and domain for a geometric sequence with a first term of 2 and a second term of -8?

OpenStudy (anonymous):

a sub(n)= 2(-8)^n-1; all integers where n is greater than or equal to 1 a sub(n)= 2(-8)^n-1; all integers where n is greater than or equal to 0 a sub(n)= 2(-4)^n-1; all integers where n is greater than or equal to 0 a sub(n)= 2(-4)^n-1; all integers where n is greater than or equal to 1

OpenStudy (cwrw238):

general form is ar^(n-1) where a = first term and r = common ratio, n = term number common ratio may be found by dividing second term by the first term

OpenStudy (anonymous):

-4=r a=2 n=n

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

now you need to look at values of n to find the domain

OpenStudy (anonymous):

so the equation s 2(-4)^n-1; all integers where n is greater than or equal to ??

OpenStudy (cwrw238):

right so far - the first term is 2 what value must n be for first term to be 2?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

Oops! 1

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

the second time?

OpenStudy (cwrw238):

so the last option is correct one

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

2* (-4)^(1-1) = 2* (-4)^0 = 2*1 = 2

OpenStudy (anonymous):

Thank you!!